### v1

# Bounds and Bugs: The Limits of Symmetry Metrics to Detect Partisan Gerrymandering

Abstract
We provide both a theoretical and empirical analysis of the Mean-Median Difference (MM) and Partisan Bias (PB), which are both symmetry metrics intended to detect gerrymandering. We consider vote-share, seat-share pairs $(V,S)$ for which one can construct election data having vote share $V$ and seat share $S$, and turnout is equal in each district. We calculate the range of values that MM and PB can achieve on that constructed election data. In the process, we find the range of vote-share, seat share pairs $(V,S)$ for which there is constructed election data with vote share $V$, seat share $S$, and $MM=0$, and see that the corresponding range for PB is the same set of $(V,S)$ pairs. We show how the set of such $(V,S)$ pairs allowing for $MM=0$ (and $PB=0$) changes when turnout in each district is allowed to be different.
Although the set of $(V,S)$ pairs for which there is election data with $MM=0$ is the same as the set of $(V,S)$ pairs for which there is election data with $PB=0$, the range of possible values for MM and PB on a fixed $(V,S)$ is different. Additionally, for a fixed constructed election outcome, the values of the Mean-Median Difference and Partisan Bias can theoretically be as large as 0.5. We show empirically that these two metric values can differ by as much as 0.33 in US congressional map data. We use both neutral ensemble analysis and the short-burst method to show that neither the Mean-Median Difference nor the Partisan Bias can reliably detect when a districting map has an extreme number of districts won by a particular party. Finally, we give additional empirical and logical arguments in an attempt to explain why other metrics are better at detecting when a districting map has an extreme number of districts won by a particular party.
1 Introduction
In this article, we study two partisan symmetry metrics that have been part of the academic study of gerrymandering for many decades, as well as on-the-ground expert testimony for many years: the Mean-Median Difference and the Partisan Bias. These metrics are embedded in publicly available websites to evaluate maps such as Dave’s Redistricting App [12] and PlanScore [15] , software such as Maptitude for Redistricting [5], and ensemble-creating python libraries such as Gerrychain [28]. While some excellent studies have revealed the “problems and paradoxes” of these kinds of metrics [10], they remain ubiquitous in software, libraries, and expert testimonies on partisan gerrymandering.
Here we conduct two different studies of these metrics. We first conduct a theoretical study in which, for fixed vote-share $V$ and seat share $S$, we calculate the possible range of values for the Mean-Median (MM) and Partisan Bias (PB) on constructed election data having vote share $V$ and seat share $S$ when turnout is equal in each district. As a corollary, we prove which vote-share, seat-share pairs $(V,S)$ have some constructed corresponding election outcome data with Mean-Median (MM) value 0. (A value of 0 for these metrics is a “perfect score,” indicating no gerrymandering). We do the same for Partisan Bias (PB), showing that, while the range of values for MM and PB are different for fixed $(V,S)$, the $(V,S)$ pairs allowing for $MM=0$ is the same as the $(V,S)$ pairs allowing for $PB=0$. We also show how those vote-share seat-share pairs change as voter turnout between districts changes.
Secondly, we conduct a series of empirical studies on the Mean-Median Difference and Partisan Bias. Although the set of $(V,S)$ pairs allowing for $MM=0$ is the same as the set of $(V,S)$ pairs allowing for $PB=0$, the values of the Mean-Median Difference and Partisan Bias can be different for a single election outcome. Indeed, the maximum theoretical difference between the two is 0.5, and a difference of up to 0.33 is seen in our empirical results, using current US districting maps.
We also empirically study the efficacy of these metrics through two techniques: neutral ensemble analysis, and a short burst analysis. Neutral ensemble analysis is a strategy used by many groups and researchers (see [20] and [21], for example) to detect gerrymandering. It involves creating a large number of valid redistricting maps (a neutral ensemble of maps, not chosen to optimize any value), and comparing a proposed map to the maps in the neutral ensemble. If the proposed map appears to be an outlier within the neutral ensemble, this may indicate the presence of gerrymandering. We show that both the Mean-Median Difference [^1] This is a replication study for the Mean-Median Difference, as it was first shown in [30]. The results for Partisan Bias, however, are new. and the Partisan Bias are not as effective at detecting maps with an extreme number of districts won by some party as other metrics when used on a neutral ensemble.
The Short Burst method was developed by Cannon et al [8] to create “extreme maps” (in their article, they focus on maps with extreme numbers of majority-minority districts). We use their method to produce maps with extreme numbers of Republican-won districts, and maps with extreme numbers of Democratic-won districts. We use the maps created through the Short Burst method to show that for both the Mean-Median Difference [^2] Again, this is a replication study for the Mean-Median Difference, as it was first shown in [30], and the results for Partisan Bias are new. and Partisan Bias, extreme values of those metrics do not correspond to maps with an extreme number of Republican-won districts (or an extreme number of Democratic-won districts).
1.1 Background and Literature Review
Partisan gerrymandering has recently become a rich topic of study for mathematicians and computer scientists, and the most straightforward and basic tools that have been proposed to detect gerrymandering are metrics. Historically, shape metrics such as the Polsby-Popper, Reock, and Convex Hull measures have been used to measure the irregularity of a district’s shape (see, for example, [1]). But recent studies have shown that having reasonable shape metric scores does not preclude a map from being an extreme partisan gerrymander (see, for example, [22]), and that even calculating shape metric values is subject to manipulation and abuse [4].
In the 1970s and for decades after, measures of partisan symmetry such as the Mean-Median Difference and Partisan Bias arose in the academic and legal communities [29, 16, 13, 23, 18, 26]. Symmetry metrics hinge on the idea that a map that gives party $A$ seat share $S$ when the vote share is $V$ must also give party $B$ seat share $S$ (or approximately $S$) when party $B$’s vote share is $V$ in order to not be a gerrymander. Other metrics such as the GEO metric [6], Efficiency Gap [31], and Declination [36] have been proposed in more recent years.
Of course, each of these metrics has their own shortcomings. In [35], Warrington compared the Partisan Bias, Mean-Median Difference, Efficiency Gap, and Declination on 12 hypothetical election outcomes, arguing that his Declination had fewer false positives and false negatives on those 12 constructed examples. In [10], DeFord et al showed that partisan symmetry metrics such as the Mean-Median Difference and Partisan Bias can be unable to distinguish between Republican and Democratic favoring maps on a couple of sample states, and argue that, “For those who do want to constrain the most extreme partisan outcomes that line drawing can secure, these investigations should serve as a strong caution regarding the use of partisan symmetry metrics.” Veomett [34] showed that the Efficiency Gap does not satisfy the Efficiency Principle (as its creators claimed [32]). Campisi at all [7] showed that the Declination does not catch all “packing and cracking” (the techniques understood to implement a gerrymander) as its creator claimed [36]. And in [30], Ratliff et al compared the Partisan Bias, Efficiency Gap, Declination, and GEO metrics in terms of their gameability. They argued that no metric is meaningfully different from measuring “districts won by party $A$” when paired with a neutral ensemble analysis, and that the Mean-Median difference had particularly unreliable and misleading results.
In this article, we fill a significant knowledge gap for the Mean-Median Difference and Partisan Bias metrics. Specifically, given a metric $m$, a very natural question is: For which vote-share, seat share pairs $(V,S)$ do there exist election data with vote share $V$, seat share $S$, and metric $m$ indicates no gerrymandering? This question has been answered for the Efficiency Gap [34] and the Declination [7]. Here, we answer that previously unanswered question for the Mean-Median Difference and the Partisan Bias. Indeed, we go even further by showing the potential range of values for MM and PB for a fixed vote-share, seat-share pair $(V,S)$.
We will also expand the empirical studies on symmetry metrics mentioned above. The authors of [10] considered three states, with a single election outcome used in the partisan analysis for each state. The authors of [30] used 18 different maps (3 maps for each of 6 states), but a single election outcome for each state. Here, we use the same maps considered in [30] but consider three elections per state in our neutral ensemble analysis study. We also add the partisan bias, which was not considered in [30]. Through these additional studies, we affirm the conclusion in [10] quoted above (that the Mean-Median Difference and Partisan Bias cannot constrain extreme partisan outcomes), and show that Partisan Bias is as erratic as the Mean-Median Difference was shown to be in [30].
We add additional empirical studies showing how MM and PB can differ in real-world data. We also provide logical and empirical arguments for why the Efficiency Gap, Declination, and GEO metric track “districts won” better than MM and PB.
1.2 “Districts Won” as the Foundation of Gerrymandering
The ultimate goal of a gerrymandering mapmaker is focused on the districts that are won (or projected to be won) by some party or group. It is for this reason that we focus our neutral ensemble analysis and short burst analysis on a metric’s ability to indicate when a party is winning an extreme number of districts in a state.
It may be the case that a metric is able to detect an extreme number of districts won, and also detect something additional about a map. For example, we will see that the GEO metric is effective at detecting extreme maps, and others have argued that the GEO metric is able to give additional information about which districts or regions of the state are potentially gerrymandered [6].
It may also be the case that a metric is unable to detect an extreme number of districts won, but detects something else entirely. We argue that this is the case for the Mean-Median Difference and Partisan Bias. As described in Section 1.3, Section 1.4, and [10], these two metrics are symmetry metrics that can show the asymmetry in outcomes for the two parties under an assumption of uniform partisan swing. However, as we shall see, these metrics are not reliable at detecting an extreme number of districts won. Since the districts won is the focus of a gerrymandering mapmaker, we argue that this implies that MM and PB are not effective at detecting gerrymandering.
1.3 Definitions
The metrics under consideration in this article are the Mean-Median Difference and the Partisan Bias. These metrics are well-known in the community (see, for example, [15]), but for completeness we give their definitions here. We shall also show how they can be visualized on the seats-votes curve.
Definition 1.
Consider an election with $n$ districts. Let $V_{1}\leq V_{2}\leq\cdots\leq V_{n}$ be the vote shares for party $A$. Then the Mean-Median Difference for this election is
$MM=\text{median}\left\{V_{1},V_{2},\dots,V_{n}\right\}-\text{mean}\left\{V_{1},V_{2},\dots,V_{n}\right\}$
Definition 2.
Consider an election with $n$ districts. Let $V_{1}\leq V_{2}\leq\cdots\leq V_{n}$ be the vote shares for party $A$, and let $\overline{V}=\text{mean}\left\{V_{1},V_{2},\dots,V_{n}\right\}$ Then the Partisan Bias for this election is
$PB=\frac{1}{2}\left(\text{Proportion of }V_{i}\text{ larger than }\overline{V}-\text{Proportion of }V_{i}\text{ smaller than }\overline{V}\right)$
In the next couple of sections, (Sections 1.4 and 1.5) we assume that the statewide vote share $V$ is simply the average of the vote shares in each district: $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. In particular, note that the creation of the seats-votes curve assumes that the statewide vote share $V$ satisfies $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. This is true when turnout in each district is the same (though is not true in general). We drop this assumption in Section 1.6.
1.4 The Seats-Votes Curve
Both the Mean-Median Difference and the Partisan Bias can be visualized on the seats-votes curve corresponding to the election with data $V_{1},V_{2},\dots,V_{n}$. The seats-votes curve is a curve drawn in the vote-share/seat-share plane under the assumption of uniform partisan swing. More specifically, the data for a single election corresponds to a vote share $V$ (for now assumed to be the mean of the $V_{i}$s), and a seat share $S$. One assumes that the vote shares $V_{i}$ swing towards party $A$ in a uniform manner; they all increase by the same vote share or decrease by the same vote share. This results in a new vote share $V^{\prime}$, and (depending on whether that vote share swing was sufficient to flip a district) a potentially new seat share $S^{\prime}$. All such new $(V^{\prime},S^{\prime})$ are drawn in the plane, which results in a curve which is a step function. Two examples can be seen in Figure 1 below. One curve corresponds to the seats-votes curve for a map whose vote shares are 0.2, 0.3, 0.55, 0.6, and 0.65; the other corresponds to real data from the Massachusetts 2011 congressional map.
Figure 1: On the left: seats-votes curve for an election with districts receiving vote shares of 0.2, 0.3, 0.55, 0.6, and 0.65. On the right: seats-votes curve for Massachusetts 2011 congressional map and the 2016 presidential election. Massachusetts data from [19].
It is well-known (see, for example, [10]) that the Mean-Median difference is the distance between the point (0.5, 0.5) and the intersection point between the seats-votes curve and the line $S=0.5$. Indeed, one can see from the figure that the distance between the point (0.5, 0.5) and the intersection of the seats-votes curve with the line $S=0.5$ is a little under 0.1. If we calculate its value using Definition 1, we have $MM=0.09$. Because of the way it can be visualized on the seats-votes curve, the Mean-Median Difference is described as the amount the vote share can fall below (or exceed) 50% while resulting in a seat share of 50%.
It is also well-known (see again [10]) that the Partisan Bias is the distance between the point (0.5, 0.5) and the intersection point between the seats-votes curve and the line $V=0.5$. One can see from the figure that the distance between the point (0.5, 0.5) and the intersection of the seats-votes curve with the line $V=0.5$ is approximately 0.1. If we calculate its value using Definition 2, we have $PB=0.1$. Because of the way it can be visualized on the seats-votes curve, the Partisan Bias is described as the amount the seat share can exceed (or fall below) 50% when the vote share is 50%.
Please note that the correspondence between the $MM$, $PB$, and the seats-votes curve described above is only true when one can reach $S=0$ and $S=1$ under uniform partisan swing. That is, this is assuming that in order to reach $S=1$ (or $S=0$) under uniform partisan swing, we don’t push a district’s vote share above 1 (or below 0) [^3] If $V_{1}$ is party $A$’s minimum vote share in a district and $V_{n}$ is party $A$’s maximum vote share in a district, then we can mathematically state this as follows: $0.5-V_{1}<1-V_{n}$ and $V_{n}-0.5<V_{1}$.
Of course, uniform partisan swing is by no means guaranteed or even likely to happen in practice (see [37] for other models of swing). Nevertheless, metrics derived from the seats-votes curve, referred to as partisan symmetry metrics, were the first election outcome metrics introduced to detect partisan gerrymandering, and have had significant staying power (as mentioned in Section 1). Indeed, the authors of [10] pointed out amicus briefs that have used partisan symmetry metrics, including LULAC v. Perry (2006) [24], Whitford v. Gill (2018) [14], and Rucho v. Common Cause (2019) [17]. Partisan symmetry metrics have also been used in various expert testimony reports, including Session v. Perry (2003) [2], Perez v Perry (2011) [27], Favors v. Cuomo (2012) [3], and an appeal of Clarke v. Wisconsin Elections Commission (2023) [9]. We hope that with articles such as [10], [30], and this article, experts will use symmetry metrics such as the Mean-Median Difference and Partisan Bias with significantly more caution.
1.5 Relationship between Mean-Median Difference and Partisan Bias
The Mean-Median Difference and Partisan Bias metrics cannot disagree in sign: if one is positive, the other is non-negative, and if one is negative the other is not positive. This follows directly from the fact that the seats-votes curve is a non-decreasing curve going through points (0,0) and (1,1). Note that whenever $MM=0$, the median of the $V_{i}$s is equal to the mean. This implies that the number of $V_{i}$ above the mean must be equal to the number of $V_{i}$s below; in other words, it implies that $PB=0$.
However, it is possible for $PB=0$ without $MM=0$ for a specific election. For example, if an election had 4 districts with vote shares 0.55, 0.6, 0.7, and 0.9, $\overline{V}=0.6875$. Since the number of districts with vote shares larger than $\overline{V}$ is the same as the number of districts with vote shares less than $\overline{V}$, $PB=0$. However, the median vote share is 0.65, implying $MM=0.65-0.6875\not=0$.
Nevertheless, as we shall see in Section 2, the vote-share, seat-share pairs $(V,S)$ for which there exists constructed election data with vote share $V$, seat share $S$, and $PB=0$ is the same as the vote-share seat-share pairs corresponding to constructed election data with $MM=0$. When one considers the fact that seat shares $S=\frac{k}{n}$ can correspond to maps with an arbitrarily large number of districts (by scaling both the numerator and denominator by some large $M$), this fact is not difficult to believe (at least in the limit). For example, suppose $V_{1},V_{2},\dots,V_{n}$ are district vote shares corresponding to statewide vote share $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$, seat share $S$, and $PB=0$. We can then consider an election with $Mn+2$ districts, having $M$ districts of vote share $V_{i}$ for $i=1,2,\dots,n$, and two districts of vote share $V$. The mean of these districts’ vote shares will still be $V$, and the seat share will be very close to $S$ (the larger $M$ is, the closer to $S$ the seat share will be). And because the original districting map had an equal number of districts with vote share above $V$ as it had below, the median of this new election will be $V$, implying that $MM=0$ for this new election.
The supremal difference between $PB$ and $MM$ is 0.5. This can be argued algebraically, but it can be seen more easily by considering how the $MM$ and $PB$ correspond to points on the seats-votes curve. For example, if $MM<0$, then the supremal value of $MM-PB$ occurs when $PB$ is near -0.5 and $MM$ is near 0. An example of such a seats-votes curve can be seen in Figure 2.
Figure 2: A Seats-Votes curve whose value of $MM-PB$ is near the maximal value of 0.5. This curve is the result of an election with 19 districts having vote share 0.49 and one district having vote share 0.89.
1.6 Unequal Turnout
Of course, no state has identical turnout in every district, so if the statewide vote share is $V$ and the district vote shares are $V_{1},V_{2},\dots,V_{n}$, then in general, $V\not=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. Indeed, $V=\sum_{i=1}^{n}\alpha_{i}V_{i}$ where $\sum_{i=1}^{n}\alpha_{i}=1$, and the non-negative weights $\alpha_{i}$ are given by turnout:
$\alpha_{i}=\frac{\text{turnout in district }i}{\text{turnout throughout the state}}$
As we shall see in Section 2, if we consider election data corresponding to uneven turnout between districts, we can get a wider array of $(V,S)$ pairs giving $MM=0$ and $PB=0$.
2 Theoretical Results
We first note that, if an election has vote-share $V$ and seat-share $S$, then we must have $0\leq V\leq 1,0\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$ (see, for example, [33]).
Please note that many of our constructed election data cannot be visualized on the seats-votes curve under the assumption of uniform partisan swing. (Recall from Section 1.4 that this may be the case if, under uniform partisan swing, we must have some districts with vote share less than 0 in order to achieve $S=0$ and/or districts with vote share more than 1 in order to achieve $S=1$). Also please note that our constructed data does not limit the number of districts. The main theoretical results of this article are below.
Theorem 1.
Suppose $(V,S)$ is a pair of rational numbers with $0\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$. Then there exists constructed election data with vote share $V$, seat share $S$, turnout is equal in all districts, and Partisan Bias $PB$ for any $PB$ in the following ranges:
[…]
Note that if party $A$’s seat share is $S$, then $1-S$ is party $B$’s seat share, and that party $A$’s PB value (and MM value) is the negation of party $B$’s PB value (MM value). Thus, Theorem 1 gives all possible range of values for PB for any possible $(V,S)$. These minimum and maximum values can be visualized in Figure 3. Theorem 1 is proved in Appendix A.
Figure 3: Range of values for Partisan Bias (using constructed election data) for each vote-share, seat-share pair $(V,S)$, when turnout in each district is equal. The hue at $(V,S)$ in the image on the left corresponds to the minimum possible value of PB. The hue at $(V,S)$ in the image on the right corresponds to the maximum possible value of PB. All values in between are also achievable on some constructed election data.
In Figure 3, we can visually see a discontinuity at $V=\frac{1}{2}$. We can understand this intuitively as well, by using the following general argument: fix $V\not=\frac{1}{2}$ and $S\geq\frac{1}{2}$. Then we can construct election data with the correct number of winning and losing districts such that the number of districts whose vote share is equal to $V$ is maximized. If we barely raise the vote shares of nearly all of the districts whose vote share is $V$ (and lower a single seat share), this maximizes PB, and keeps the correct vote share and seat share. If we barely lower the vote shares of nearly all of the districts whose vote share is $V$ (and raise a single seat share), this minimizes PB and keeps the correct vote share and seat share.
However, this argument breaks down when $V=\frac{1}{2}$ because barely raising or lowering the vote share can change it from winning to losing or vice versa. This creates the discontinuity at that line.
We also have the corresponding Theorem for the Mean-Median Difference:
Theorem 2.
Suppose $(V,S)$ is a pair of rational numbers with $0\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$. Then there exists constructed election data with vote share $V$, seat share $S$, turnout is equal in all districts, and Mean-Median Difference $MM$ for any $MM$ in the following ranges:
[…]
These minimum and maximum values can be visualized in Figure 4. Theorem 2 is also proved in Appendix A.
Figure 4: Range of values for Mean-Median Difference (using constructed election data) for each vote-share, seat-share pair $(V,S)$, when turnout in each district is equal. The hue at $(V,S)$ in the image on the left corresponds to the minimum possible value of MM. The hue at $(V,S)$ in the image on the right corresponds to the maximum possible value of MM. All values in between are also achievable on some constructed election data.
In Figure 4, we can visually see a discontinuity at $S=\frac{1}{2}$. This discontinuity can be understood intuitively as follows: fix $\frac{1}{2}\leq V\leq\frac{3}{4}$. When $S>\frac{1}{2}$ gets close enough to $\frac{1}{2}$, the largest that the median could possibly be is 1 and smallest the median could possibly be is $\frac{1}{2}$. Whereas when $S<\frac{1}{2}$ the largest the median could possibly be is $\frac{1}{2}$ and smallest it could possibly be is 0. However, the mean is fixed at $V$. This creates the discontinuity at $S=\frac{1}{2}$.
From Figures 3 and 4, we can see that the range of values that MM and PB can take respond differently to changing $V$ and $S$. Consider, for example, the feasible region with $\frac{1}{2}\leq S\leq 1$, $\frac{1}{2}\leq V\leq 1$. For a fixed $S$ in that region, the maximum possible value for $PB$ doesn’t depend on $V$ at all, as it is $S-\frac{1}{2}$. Additionally, note that for $V$ just barely less than $\frac{1}{2}$, PB can achieve nearly the maximal value of $\frac{1}{2}$, regardless of $S$. (And symmetrically, for $V$ just barely above $\frac{1}{2}$, PB can achieve nearly the minimal value of $-\frac{1}{2}$).
For the Mean-Median Difference, we can consider the feasible region with $\frac{1}{2}\leq S\leq 1$, $\frac{1}{4}\leq V\leq\frac{3}{4}$. In this region, for a fixed $V$, regardless of the value of $S$, the Mean-Median Difference has the same possible minimum value ($\frac{1}{2}-V$), while the largest maximum values occur near the line $S=2V-\frac{1}{2}$, which happens to be the line where the Efficiency Gap is always 0. Indeed, the Efficiency Gap can only take on a single value when turnout in each district is equal, which can be seen in Figure 5.
Figure 5: Only possible value of Efficiency Gap (using constructed election data) for each vote-share, seat-share pair $(V,S)$, when turnout in each district is equal.
While these observations may be mathematically interesting, some of them seem potentially problematic in a metric intended to detect partisan gerrymandering. For example, when considering the Partisan Bias: when $S$ with $\frac{1}{2}\leq V\leq 1$ is fixed, but $V$ increases, why should PB be able to take on values indicating the same amount of benefit to the party in question? And why should nearly the maximum benefit be able to be indicated for $V$ just under $\frac{1}{2}$ and $S$ any value? For the Mean-Median Difference: for fixed $V$, why would the largest possible values be achievable on the line $S=2V-\frac{1}{2}$?
These questions are unanswerable, as these mathematical effects are surely not what was considered in the design of the Mean-Median Difference and Partisan Bias. Nevertheless, we pose them simply to point out the likely unintended consequences of the definitions of MM and PB.
Note from Theorems 1 and 2, that when $S\geq\frac{1}{2}$, both the Mean-Median Difference and Partisan Bias cannot be zero for $V<\frac{1}{2}$, and that both the Mean-Median Difference and Partisan Bias can be zero for any $V$ with $\frac{1}{2}\leq V<\frac{3}{4}$. We can deduce the remainder of Corollary 1 by setting the lower bound equal to 0 in the region $\frac{3}{4}\leq V\leq 1$.
Corollary 1.
Suppose $V$ is a rational number with $\frac{1}{2}\leq V<\frac{3}{4}$ and $S$ is any rational number with $\frac{1}{2}\leq S\leq 1$. Alternatively, suppose that $V$ is a rational number with $\frac{3}{4}\leq V\leq 1$ and $S$ is any rational number with $\frac{3V-2}{2V-1}<S\leq 1$. Then one can construct election data with seat share $S$, vote share $V$, turnout in each district is equal, and $PB=0$. Additionally, one can construct election data with seat share $S$, vote share $V$, turnout in each district is equal, and $MM=0$.
Note that, by symmetry, we have all $(V,S)$ pairs for which there exists constructed election data with vote share $V$, seat share $S$, and Partisan Bias 0. This result shows that, when turnout in each district is equal, a very wide array of $(V,S)$ pairs can result in $MM=0$ and $PB=0$. The region consisting of all $(V,S)$ for which there is constructed election data with $MM=0$ and $PB=0$ can be seen in Figure 6.
Figure 6: Region of all $(V,S)$ for which there is constructed election data with turnout in all districts equal and $MM=0$ (which is the same as the region of all $(V,S)$ for which there is constructed election data with turnout in all districts equal and $PB=0$).
This is different from the set of $(V,S)$ for which there is constructed election data with turnout equal in all districts and $\text{Declination}=0$ [7] (see Figure 7), and also quite different from the set of $(V,S)$ for which there is constructed election data with turnout equal in all districts and $EG=0$ [34] (see Figure 8). From these Figures, we can deduce that the Partisan Bias and Mean-Median Difference are the least restrictive measures in terms of which $(V,S)$ pairs are allowed to potentially have a metric value which indicates that no gerrymandering has occurred.
Figure 7: Region of all $(V,S)$ for which there is constructed election data with turnout in all districts equal and $\text{Declination}=0$ (dark grey), overlayed with the region of all $(V,S)$ for which there is constructed election data with turnout in all districts equal and $PB/MM=0$ (light grey).
Figure 8: Region of all $(V,S)$ for which there is constructed election data with turnout in all districts equal and $EG=0$ (dark grey), overlayed with the region of all $(V,S)$ for which there is election data with turnout in all districts equal and $PB/MM=0$ (light grey).
Note that Corollary 1 assumes that turnout in each district is equal, which is never the case in real elections. This raises the question: if turnout is unequal, say the maximum turnout divided by minimum turnout is $C$, what $(V,S)$ pairs allow for Partisan Bias and Mean-Median Difference 0? The following result answers this question.
Theorem 3.
Fix the maximum turnout to minimum turnout ratio at $C\geq 1$. Suppose that $S\geq\frac{1}{2}$ for rational seat share $S$. Then for any rational vote share $V^{*}$ with
$\frac{1}{2(S+C(1-S))}\leq V^{*}<\frac{1+C(3-2S)}{(C+1)(3-2S)}$
there is constructed election data with vote share $V^{*}$, seat share $S$, and $PB=0$. Additionally, for any such $V^{*}$ there is constructed election data with vote share $V^{*}$, seat share $S$, with $MM=0$.
Note that Theorem 3 gives the lower bound of $\frac{1}{2}$ when $C=1$ (when turnout in all districts is equal), which is what we already know to be true from Corollary 1. For the upper bound, when $C=1$ we again get the result of Corollary 1. Specifically, setting $C=1$, Theorem 3 gives $V^{*}\leq\frac{2-S}{3-2S}$, which is the result of solving the equation $S=\frac{3V^{*}-2}{2V^{*}-1}$ (from Corollary 1) for $V^{*}$. When turnout is unequal, Theorem 3 expands the pairs $(V,S)$ for which there is election data with vote share $V$, seat share $S$, and Partisan Bias/Mean-Median Difference values of 0. For example, Figure 9 shows how this region expands for different values of $C$. As $C\to\infty$, the region expands to fill the $[0,1]\times[0,1]$ square.
Figure 9: Region of all $(V,S)$ for which there is constructed election data with $PB/MM=0$. The dark grey portion corresponds to turnout ratio $C=1$ and the light grey portion corresponds to turnout ratio $C=4$ in the left image, and $C=100$ in the right image.
We note that in [34], the author evaluated election turnout ratios for all states with at least eight congressional districts, using data from 2014 and 2016, and showed that the ratio of maximum turnout to minimum turnout can be as high as 4.4 (in California). The 2022 general congressional district elections in California had a turnout ratio of 3.29, suggesting that high turnout ratios remain something that is seen in practice, even in general elections.
Theorem 3 is also proven in Section A.
3 Empirical Results: Difference between the Mean-Median and Partisan Bias
From Section 2, we see that the set of $(V,S)$ pairs for which there is election data giving $MM=0$ is the same as the set of $(V,S)$ pairs for which there is election data giving $PB=0$. However, the actual values of the Mean-Median Difference and Partisan Bias can be different on a single election. In this section, we empirically study the $(MM,PB)$ pairs found in many US maps.
All of the data for this study were collected from Dave’s Redistricting App [12] between April 19 and April 21, 2024. For each state, all districting maps available were collected; this typically consisted of Congressional districts, State House districts, and State Senate districts [^4] Several state house and state senate districting maps have districts with extremely different populations, for various reasons. For example, the Vermont state senate has each district being a multimember district, but with a range of 1-3 members per district. Those maps were simply included as-is, and did not give the most extreme values of $MM-PB$ that we observed., with exceptions for states that have one congressional district (consisting of the whole state) and/or a single state legislative map. (If a state has a single state legislative map, that map was included in the “State Senate” maps). The maps collected were those listed as “current maps” according to DRA, and using the partisan data on the site (which is a composite of several election results).
We can see a scatter plot of $(MM,PB)$ pairs for all of the maps analyzed in Figure 10; there are 44 Congressional maps, 45 State House maps, and 50 State Senate maps for a total of 139 maps all together. Positive values are intended to indicate a map that benefits the Democratic party; negative values are intended to indicate a map that benefits the Republican party.
From these scatterplots, we can see that the Mean-Median Difference and Partisan Bias values can be quite different in real-world data. Recall from Section 1.5 that the supremal theoretical difference between the Mean-Median Difference and Partisan Bias is 0.5. The largest difference seen in our data was 0.33, which was achieved on South Carolina’s congressional map. Other maps with a large difference (in absolute value) were Massachusetts congressional (0.25), Indiana Congressional (0.23), Wisconsin Congressional (0.20), Iowa Congressional (-0.22), and Nevada Congressional (-0.22). The only states having both Mean-Median Difference and Partisan Bias values of 0 were those where it is guaranteed to have $MM=PB=0$: states with two congressional districts (Hawaii, Idaho, Maine, Montana, New Hampshire, Rhode Island, and West Virginia).
Recall that values far from 0 for these metrics is intended to indicate a greater degree of gerrymandering, while the sign is intended to indicate which party was gerrymandered. As stated in Section 1.5, the sign of the two metrics cannot differ, although the values can. These results indicate that the Mean-Median Difference and Partisan Bias metrics can disagree quite drastically as to the degree of gerrymandering happening within a state.
We also note that the sign of these metrics (positive indicating a map benefitting Democrats and negative indicating a map benefitting Republicans) on certain states may be surprising to some readers. For example, the Partisan Bias and Mean-Median Difference suggest that the Massachusetts congressional map is potentially a Republican gerrymander. The seats-votes curve for that map was in Figure 1; we encourage the reader to explore these plots and consider how frequently the sign of these partisan symmetry metrics are incorrect (the incorrectness of sign was also addressed in [10]).
Figure 10: $(MM,PB)$ pairs for all for State House, State Senate, and Congressional districts, all states. Data gathered from Dave’s Redistricting App between April 19 and April 21, 2024, for maps listed as “current maps” at the time. Top image allows points to be more easily distinguished, bottom image has the same scale on the vertical and horizontal axes. Dotted lines correspond to $MM=PB$.
4 Empirical Results: Accuracy in a Neutral Ensemble and Reliability
From Section 2, we see that both the Mean-Median Difference and Partisan Bias allow for a wide variety of seats/votes pairs corresponding to election data with metric value of 0. This suggests that potentially these metrics are less powerful at detecting gerrymanders. Here we provide empirical results that support this suggestion.
4.1 Outlier Analysis
One of the techniques of detecting gerrymandering that is most widely accepted by the mathematical and computational community is outlier analysis. Through this technique, a neutral ensemble of typically tens of thousands of redistricting maps is created, to which the proposed map is compared. If the proposed map looks like an outlier compared to the other maps in the ensemble, this could suggest the presence of gerrymandering. Several research groups implement different varieties of this outlier analysis; we refer the reader to the Metric Geometry Gerrymandering group [20] or the Quantifying Gerrymandering group [21] for examples.
This outlier analysis can use any kind of metric, from the Mean-Median Difference or Partisan Bias, GEO metric, Efficiency Gap, to the simple metric of number of districts won by a particular party. It is a natural question to ask: does the Mean-Median Difference or Partisan Bias seem to do a good job when it comes to neutral ensembles? That is, does it agree with the other metrics in determining which maps are gerrymanders? Or does it have positive errors (detecting a gerrymander where there is none) or negative errors (missing a clear gerrymander)? Note that we consider the number of districts won to be the foundation of a partisan gerrymander (see the discussion in Section 1.2).
This question was answered for the Mean-Median Difference in [30], which showed that the Mean-Median Difference commits both positive and negative errors. Here we show the same for the Partisan Bias.
For the purposes of this study, we used the exact same 18 maps that were studied by the authors of [30], but use three elections for each state ([30] considered one election per state). Those 18 maps corresponded to 3 maps per state in 6 states which have different political geographies: Massachusetts, Michigan, Oklahoma, Oregon, Pennsylvania, and Texas. Data for each state was taken from the Metric Geometry Gerrymandering Group’s GitHub repository containing data from the maps created after the 2010 Census cycle [19]. For each state, we used the most recent Gubenatorial, Presidential, and US Senate election that was available in the datset at [19], with the exception of Oklahoma, which had no Presidential or US Senate election data (we chose the 2018 Attorney General and 2018 Lieutenant Governor elections instead). We also use the same Python library of gerrychain [28] that was used in [30]; gerrychian implements a Markov Chain Monte Carlo process in order to create the neutral ensemble of potential redistricting maps. We used a population deviation of 5% for congressional maps and 11% for state house/state senate maps in all of our ensembles, and we created an ensemble of size 20,000 (also mimicking what was used in [30]).
A sample result can be seen in Figure 11. Please note that the GEO metric is somewhat different from other metrics in that large values of a Democratic GEO score indicate a map that is less beneficial to the Democratic party, whereas for all the other metrics a large value indicates a map that is more beneficial to the Democratic party. For this reason, we display the negated Democratic GEO score in our histograms, so that for all histograms smaller values are less beneficial to the Democratic party. We also note that the creators of the GEO score emphasized that both the Democratic and Republican GEO scores should be used together (not separately) to evaluate a map [6].
Figure 11: Neutral ensembles for Massachusett’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Here we can see that all of the metrics with all elections considered suggest the same thing: the enacted map looks fairly typical compared to the maps in the neutral ensemble. However, three of the 18 considered maps revealed that the Partisan Bias, like the Mean-Median Difference, commits both positive and negative errors. Consider Figures 12 and 13.
Figure 12: Neutral ensembles for Massachusett’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric. Declination not included because it is not defined for nearly all maps. Partisan Bias suggests that this map is a gerrymander on all election data, unlike all other metrics.
Figure 13: Neutral ensembles for Texas’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric. Partisan Bias suggests that this map is a gerrymander on all election data, unlike all other metrics. Both GEO scores paired together raise a possibility of gerrymandering on the GOV 14 data.
In Figures 12 and 13 we can see that both the Massachusetts and Texas Congressional districting maps are flagged as outliers by the Partisan Bias metric on all election data, unlike any other metric. Additionally, research has clearly shown that, while MA has all 9 districts going to the Democratic party, it is essentially impossible to draw a map in which Republicans can win a single district [11] (indicating that the Massachusetts Congressional map is not a gerrymander).
Alternatively, we can see in Figure 14 that the Partisan Bias does not flag Pennsylvania’s 2011 congressional map as an outlier on the two election data that all other metrics do (aside from the Mean-Median Difference, which does not detect the gerrymander on the T16SEN data). This is despite the fact that the other metrics do flag it as an outlier (except for the Mean-Median Difference!) and the wide acceptance that that map was indeed a gerrymander (including by the Pennsylvania Supreme Court [25]).
Figure 14: Neutral ensembles for Pennsylvania’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric. All metrics suggest gerrymandering on T16SEN and GOV14 data, but not on T16PRES data, except for Partisan Bias (which never detects gerrymandering) and Mean-Median Difference (which does not detect it on the T16SEN data). Recall that both GEO scores must be used together.
4.2 Short Burst Analysis
The authors of [30] also used a Short Burst analysis to test the “gameability” of some metrics intended to detect gerrymandering (specifically, the GEO Metric, Mean-Median Difference, Declination, Efficiency Gap). They define a metric as “gameable” if the metric value can stay within acceptable bounds [^5] For a metric $m$, they considered “acceptable bounds” to be $-0.16\inf(m)\leq m\leq 0.15\sup(m)$, based on the bounds recommended by the creators of the Efficiency Gap[31]., while allowing a map with the most extreme number of Democratic-won (or Republican-won) districts that can be achieved by that map.
In order to find such extreme maps, they use the Short Burst method introduced by Cannon et all in [8] [^6] The idea behind short bursts was suggested by Z. Schutzman at the Voting Rights Data Institute in the summer of 2018. In their article, Cannon et al used Short Bursts to find the largest number of majority-minority districts that a state can have in a single map. The short burst method takes a small number of steps in a Markov Chain (say 10 steps). At the end of that “short burst”, the map with the highest number of majority-minority districts is then the seed for the next short burst of size 10. This process is repeated thousands of times in an effort to find the maximum number of majority-minority districts achievable. In [8] Cannon et al showed that short bursts work better than biased walks in Markov Chains in order to find that maximum. Note that the Short Burst method can be thought of as a non-deterministic version of beam search on the metagraph [^7] The “metagraph” here is the graph whose nodes are maps. Two maps are adjacent if one can be obtained by the other through a single step in the Markov Chain constructing the ensemble of maps. In this case, that is through a single ReCom move in GerryChain. of redistricting maps, using depth-first search instead of breadth-first, and in which a random selection of maps near to the current map are explored.
In [30], Ratliff et al used this same technique, but in order to maximize the number of Democratic-won or Republican-won districts. We use the same technique here, on the same 18 maps that were discussed in Section 4.1 and using the same election data used in [30] [^8] We used a single election outcome (the same used in [30]) because the goal here is not to assess whether a map is an outlier, but just to assess bounds on the metric for maps with a fixed number of districts won.. Perhaps the most startling results in [30] were that, while the GEO Metric, Declination, and Efficiency Gap values tended to go up when the number of Democratic-won districts increased (indicating a map that benefits the Democratic party), the Mean-Median Difference values did not. The authors of [30] did not evaluate the Partisan Bias (likely because that metric is less frequently discussed in recent academic research), but we do so here. We show that the Partisan Bias shares the same uninformative behavior of the Mean-Median Difference.
For example, consider Figures 15 and 16.
Figure 15: Short burst results for Massachusetts’s State House map and 2018 US Senate election data, searching for maps with as many Democratic-won districts as possible. Declination not included because it is frequently undefined. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Note that Partisan Bias and Mean-Median values do not increase as number of districts won increases, as the other metrics do.
Figure 16: Short burst results for Massachusetts’s State House map and 2018 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Declination not included because it is frequently undefined. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Note that Partisan Bias and Mean-Median values do not decrease as number of districts won increases, as the other metrics do.
Here we can see that the range of values for all metrics except MM and PB is higher when the number of Democratic-won districts is higher. And conversely, the range of values for all metrics except MM and PB is lower when the number of Republic-won districts is higher. For the Mean-Median Difference and Partisan Bias, the range of values tends to be similar, regardless of the number of districts won by a particular party. These are the results found in [30], except that [30] did not consider the Partisan Bias, which is included here. Thus, we can similarly conclude that the Partisan Bias is not useful in determining extreme maps, just like the Mean-Median Difference.
We note that these results (indicating that the symmetry measures of Partisan Bias and Mean-Median Difference values are not useful in flagging extreme maps) are similar to those in [10], although here we see those results on 18 maps, rather than the three maps studied in [10]. The results for the other maps can be found in Appendix C.
5 Discussion
The results from our Short Burst analyses, such as those in Figures 15 and 16, show that the values of the Partisan Bias and Mean-Median difference do not indicate whether or not a map is an extreme map. The values do not generally increase as the number of Democratic-won districts increase (or decrease as the number of Republican-won districts decrease) as the values of the other metrics do. This observation, along with our discussion of “districts won” being the foundation of gerrymandering outlined in Section 1.2, leads us to suggest that the Partisan Bias and Mean-Median Difference are not accurate metrics at determining partisan gerrymandering.
But these results also invite the corresponding question: Why are the other metrics more consistent with districts won? What is it about the other metrics that leads them to increase when Democratic-won districts increase, and decrease when Democratic-won districts decrease? The answer is clear with the Efficiency Gap, as it is simply a function in the vote-share, seat-share, and a turnout ratio $\rho$ [34]. Specifically, if
$\rho=\frac{\text{average turnout in districts party }A\text{ lost}}{\text{average turnout in districts party }A\text{ won}}$
then
$EG=\left(S-\frac{1}{2}\right)-2\left(V-\frac{1}{2}\right)+\frac{S(1-S)(1-\rho)}{S(1-\rho)+\rho}$
Thus, for a neutral ensemble of maps using fixed partisan data (and thus fixed $V$), as $S$ goes up, EG will also increase [^9] Yes, $\rho$ will also change, but its change is relatively small. This is why the bands of potential EG values for different $S$ are narrow..
For the Geography and Election Outcome (GEO) metric, the argument is more intuitive. Recall that the Democratic GEO score of a map is the number of additional districts that could have potentially been made competitive for the Democratic party, and the Republican GEO score is the number of additional districts that could have potentially been made competitive for the Republican party [6]. The GEO score ratio used in the Short Burst analysis (used because it has values between -1 and 1, similar to other metrics, and thus we could display the green “reasonable band” as with the other metrics) is:
Rep GEO scoreDem GEO scoretotal number of $\frac{\text{Rep GEO score}-\text{Dem GEO score}}{\text{total number of districts}}$
For a neutral ensemble of maps, evaluated using the same partisan data (and thus fixed $V$), consider two maps with different Democratic seat shares. One can imagine that the map $M_{\text{Dem higher}}$ with higher Democratic seat share can be constructed from the map $M_{\text{Dem lower}}$ with lower Democratic seat share by adjusting $M_{\text{Dem lower}}$. Using the ideas behind the GEO metric, this may be the result of changing some of the districts that could have become competitive for the Democratic party into districts won by the Democratic party. If this is the case, the GEO score for the Democratic party in $M_{\text{Dem higher}}$ would be lower. Conversely, those districts that are now won by the Democratic party in $M_{\text{Dem higher}}$ could become competitive for the Republican party, suggesting that the GEO score for the Republican party in $M_{\text{Dem higher}}$ would be higher. Thus, this suggests that the GEO ratio for $M_{\text{Dem higher}}$ would be larger than the GEO ratio for $M_{\text{Dem lower}}$, which would explain why the GEO ratio tends to increase as the Democratic-won districts increases in the neutral ensemble.
To explain why the Declination also tends to increase as Democratic-won districts increases, we can look to the proof of Theorem 1 in [36]. What the proof shows is the following: suppose $M_{\text{Dem lower}}$ and $M_{\text{Dem higher}}$ are maps with the same partisan data (and thus same $V$) such that $M_{\text{Dem lower}}$ has lower Democratic seat share than $M_{\text{Dem higher}}$. Let $\overline{V_{\ell}}(M)$ be the average Democratic vote share in districts that Democrats lots in map M and $\overline{V_{w}}(M)$ be the average Democratic vote share in districts that Democrats won in map $M$. Then if we have both
$\overline{V_{\ell}}(M_{\text{Dem higher}})$ $<\overline{V_{\ell}}(M_{\text{Dem lower}})$ (1)
$\overline{V_{w}}(M_{\text{Dem higher}})$ $<\overline{V_{w}}(M_{\text{Dem lower}})$ (2)
then
$\delta(M_{\text{Dem lower}})<\delta(M_{\text{Dem higher}})$
where $\delta$ is the Declination. That is, if the average vote shares in districts that Democrats lost tends to get lower as $S$ increases, and if the average vote shares in districts that Democrats won tends to get lower as $S$ increases, then the Declination tends to increase as $S$ increases. Intuitively, one can understand how this might happen through packing and cracking, although it certainly may not always occur.
For example, consider two different elections in a state with 5 districts. The first election results in two districts having vote shares of 30% and three districts having vote share 80%. The second election results in one district having vote share $\frac{7}{15}\approx 46.667\%$ and four districts having vote share $\frac{19}{30}\approx 63.333\%$. The vote shares and declination values of both elections are the same ($V=60\%,\delta=0$). However the seat shares are different. Thus, for this pair of elections, the declination cannot detect the seat share change. (We note that not both of (1) and (2) can be true for this pair of elections.)
However, while equations (1) and (2) may not always be true in constructed data, they may be frequently true in real data. To assess this empirically, we created histograms of $\overline{V_{\ell}}$ and $\overline{V_{w}}$ for the neutral ensembles we created for each map and each election. Typical results can be seen in Figure 17.
Figure 17: Histograms on the left are the average vote shares in districts Democrats lost, color coded by the number of districts won by Democrats. Histograms on the right are the average vote shares in districts Democrats won, color coded by the number of districts won by Democrats. First row is for Michigan congressional, second row is Texas Congressional, third row is Pennsylvania congressional. All are using election data from the 2016 Presidential election. All show that the histogram values tend to move higher as the number of districts won by Democrats decreases.
In Figure 17, we can see that the values of $\overline{V_{\ell}}$ do tend to increase as $S$ increases, and the values of $\overline{V_{w}}$ also tend to increase as $S$ increases. This gives a suggestion as to why the Declination tends to increase as $S$ increases.
6 Conclusion
In this article, for each possible $(V,S)$ pair that can be achieved with turnout equal in each district, we have calculated the possible range of values of $MM$ and $PB$ for constructed election data with vote-share $V$ and seat-share $S$. We have seen that this range of values behaves in ways that are not what one would naturally expect of a metric intended to detect partisan gerrymandering.
As a result of our calculations, we have completed a long-missing study of the Mean-Median Difference and Partisan Bias, which had been completed for other metrics: showing which seats-votes pairs $(V,S)$ have constructed election data giving $MM=0$ or $PB=0$. We have shown that these $(V,S)$ pairs are the same for both metrics, and cover a wide swath of the $[0,1]\times[0,1]$ square. This region expands when we allow the ratio $C$ of maximum turnout to minimum turnout to be larger than 1; so much so that the region fills the $[0,1]\times[0,1]$ square in the limit as $C\to\infty$. This suggests that the Mean-Median Difference and Partisan Bias metrics are potentially less reliable when detecting gerrymanders.
Indeed, we empirically show that, although the signs of the Mean-Median Difference and Partisan Bias must be the same, the values of the two metrics can be quite different. The largest theoretical difference is 0.5, and a difference of 0.33 is seen in South Carolina’s Congressional map (with several other maps having a difference of 0.2 or more in absolute value).
The study in [30] showed that the Mean-Median Difference can give misleading results even when evaluated on a neutral ensemble, and we show here that the same is true of the Partisan Bias.
The study in [30] showed that the range of values of the Mean-Median Difference is similar for both extreme maps (maps with an extreme number of districts won by a particular party) and non-extreme maps, and we show similar results here for the Partisan Bias. We discuss reasons why the Efficiency Gap, Declination, and GEO metric are better at tracking “districts won” than the Mean-Median Difference and Partisan Bias, and back up the reasoning for the Declination with empirical studies.
Warnings have been raised about symmetry metrics such as the Mean-Median Difference and Partisan Bias in [10], stating “For those who do want to constrain the most extreme partisan outcomes that line drawing can secure, these investigations should serve as a strong caution regarding the use of partisan symmetry metrics, whether in the plan adoption stage or in plan evaluation after subsequent elections have been conducted.” The additional theoretical and empirical findings in this article give even more evidence that partisan symmetry metric values fail to accurately detect when a map has an extreme number of districts won by a particular party. When used in a neutral ensemble, an outlier Mean-Median Difference of Partisan Bias may or may not correspond to an outlier in number of districts won by a given party. We similarly strongly caution against using these metrics at any stage. They simply are not reliable at flagging extreme maps, even within a neutral ensemble.
7 Acknowledgements
The author would like to thank Thomas Ratliff for sharing the code he wrote to make the beautiful images for the Short Burst analysis results, most of which can be seen in Appendix C. The author would also like to thank Moon Duchin, whose comments, questions, and suggestions have greatly improved this article.
Appendix A Proofs of Theorems 1, 2, and 3
Proof of Theorem 1.
Suppose $(V,S)$ is a pair of rational numbers with $S\geq\frac{1}{2}$, $S\leq 2V$, and $S\geq 2V-1$.
Case 1: $V=\frac{1}{2}$
Let $S=\frac{n-\ell}{n}$, and choose the denominator $n$ to be large. Choose $\epsilon>0$ small enough so that $\frac{1}{2}-(n-\ell)\epsilon>0$. Set $n-\ell$ districts to have vote share $V+\epsilon=\frac{1}{2}+\epsilon$; these are winning districts. Set one losing district to have vote share $V-(n-\ell)\epsilon=\frac{1}{2}-(n-\ell)\epsilon$, and the remaining losing districts to have vote share $V=\frac{1}{2}$. This constructed election has vote share $V$, seat share $S$, a maximum number of districts with vote share above $V$, and a minimum number of districts with vote share below $V$. The Partisan Bias is
$\frac{1}{2}\left(\frac{n-\ell}{n}-\frac{1}{n}\right)$
which has a supremal limit of $\frac{S}{2}$.
Now choose $\epsilon>0$ small enough so that $\frac{1}{2}+\ell\epsilon<1$. Set $\ell$ districts to have vote share $V-\epsilon=\frac{1}{2}-\epsilon$; these are losing districts. Set one losing district to have vote share $V+\ell\epsilon=\frac{1}{2}+\ell\epsilon$, and the remaining winning districts to have vote share $V=\frac{1}{2}$. This constructed election has vote share $V$, seat share $S$, a minimum number of districts with vote share above $V$, and a maximum number of districts with vote share below $V$. The Partisan Bias is
$\frac{1}{2}\left(\frac{1}{n}-\frac{\ell}{n}\right)$
which has an infimal limit of $\frac{S-1}{2}$. Thus, we have shown that for $V=\frac{1}{2}$, we can construct election data with PB for any
$\frac{S-1}{2}$ $<PB<\frac{S}{2}$
Case 2: $\frac{1}{4}\leq V<\frac{1}{2}$
Let $\ell$ be the number of losing districts so that $S=\frac{n-\ell}{n}$. Let $p$ be the number between 00 and 1 such that
$0\cdot p+V\left(1-p-\frac{n-\ell}{n}\right)+\frac{n-\ell}{n}\cdot\frac{1}{2}=V$
Note that this implies
$p=\frac{n-\ell}{n}\left(\frac{1}{2V}-1\right)=S\left(\frac{1}{2V}-1\right)$
By finding a common denominator, we can assume that that $p$’s denominator is $n$. Under the assumption that turnout in all districts is equal, having a proportion $p$ of districts having vote share 0, $S$ districts having vote share $\frac{1}{2}$ (considered winning districts), and $1-p-\frac{n-\ell}{n}$ districts having vote share $V$ gives the statewide vote share as $V$ and seat share $S$. It also maximizes the proportion of districts having vote share $V$.
Thus, if we want to make PB as large as possible for an election winning $S$ districts and vote share $V$, we first set $n$ (the number of districts) to be very large. Then construct election data with a proportion $p$ of districts having vote share 0, $S$ districts having vote share $\frac{1}{2}$ (considered winning districts), nearly $1-p-\frac{n-\ell}{n}$ districts being $\epsilon$-larger than $V$, and a single district less than $V$ for an extremely small $\epsilon$.
If we want to make PB as small as possible for an election winning $S$ districts and vote share arbitrarily close to $V$, we first set $n$ (the number of districts) to be very large. Then construct election data with a proportion $p$ of districts having vote share 0, $S$ districts having vote share $\frac{1}{2}$ (considered winning districts), nearly $1-p-\frac{n-\ell}{n}$ districts being $\epsilon$-smaller than $V$, and a single district larger than $V$ for an extremely small $\epsilon$.
Using these two constructions, we have for $(V,S)$ the following infimal and supremal bounds:
$\frac{1}{2}\left(\frac{n-\ell}{n}-\left(1-p-\frac{n-\ell}{n}+p\right)\right)$ $<PB<\frac{1}{2}\left(\frac{n-\ell}{n}+1-p-\frac{n-\ell}{n}-p\right)$
$\frac{1}{2}\left(2S-1\right)$ $<PB<\frac{1}{2}\left(1-2p\right)=\frac{1}{2}\left(1-2S\left(\frac{1}{2V}-1\right)\right)$
$S-\frac{1}{2}$ $<PB<\frac{1}{2}-S\left(\frac{1}{2V}-1\right)$
Case 3: $\frac{1}{2}<V\leq 1$
Let $\ell$ be the number of losing districts so that $S=\frac{n-\ell}{n}$. Let $p$ be the number between 00 and 1 such that
$\frac{\ell}{n}\cdot\frac{1}{2}+V\left(1-p-\frac{\ell}{n}\right)+p=V$
Note that this implies
$p=\frac{\frac{\ell}{n}\left(V-\frac{1}{2}\right)}{1-V}=\frac{(1-S)\left(V-\frac{1}{2}\right)}{1-V}$
By finding a common denominator, we can assume that that $p$’s denominator is $n$. Under the assumption that turnout in all districts is equal, having a proportion $p$ of districts having vote share 1, $\frac{\ell}{n}=1-S$ districts having vote share $\frac{1}{2}$ (considered losing districts), and $1-p-\frac{\ell}{n}$ districts having vote share $V$ gives the statewide vote share as $V$ and seat share $S$. It also maximizes the proportion of districts having vote share $V$.
Thus, to make PB as large as possible for an election winning $S$ districts and vote share $V$, we would first set $n$ (the number of districts) to be very large. Then construct election data with a proportion $p$ of districts having vote share 1, $\frac{\ell}{n}=1-S$ districts having vote share $\frac{1}{2}$ (considered losing districts), nearly $1-p-\frac{\ell}{n}$ districts being $\epsilon$-larger than $V$, and a single district smaller than $V$ for an extremely small $\epsilon$
If we want to make PB as small as possible for an election winning $S$ districts and vote share arbitrarily close to $V$, we would first set $n$ (the number of districts) to be very large. Then construct election data with a proportion $p$ of districts having vote share 1, $\frac{\ell}{n}=1-S$ districts having vote share $\frac{1}{2}$ (considered losing districts), nearly $1-p-\frac{\ell}{n}$ districts being $\epsilon$-smaller than $V$, and a single district larger than $V$ for an extremely small $\epsilon$.
Using these two constructions, we have for $(V,S)$ the following infimal and supremal bounds:
$\frac{1}{2}\left(p-\left(1-p-\frac{\ell}{n}+\frac{\ell}{n}\right)\right)$ $<PB<\frac{1}{2}\left(p+1-p-\frac{\ell}{n}-\frac{\ell}{n}\right)$
$\frac{1}{2}\left(2p-1\right)$ $<PB<\frac{1}{2}\left(1-2\frac{\ell}{n}\right)$
$\frac{(1-S)\left(V-\frac{1}{2}\right)}{1-V}-\frac{1}{2}$ $<PB<S-\frac{1}{2}$
Proof of Theorem 2.
Case 1: $S=\frac{1}{2}$.
Since $S=\frac{1}{2}$, the smallest that the median value could be is $\frac{1}{4}$, which would occur when the middle vote share values are 0 (a losing district) and $\frac{1}{2}$ (a winning district). For $\frac{1}{4}\leq V<\frac{1}{2}$, this median value is achievable by letting the other winning vote shares range from $\frac{1}{2}$ (when $V=\frac{1}{4}$) to 1 (for $V$ approaching $\frac{1}{2}$). This gives a minimum MM of $\frac{1}{4}-V$.
For $\frac{1}{2}\leq V\leq\frac{3}{4}$, let $p$ be a value such that
$\frac{1}{n}\left(\frac{n}{2}p+\frac{1}{2}+\left(\frac{n}{2}-1\right)\cdot 1\right)=V$
Note that this implies
$p=2\left(V+\frac{1}{2n}-\frac{1}{2}\right)\approx 2V-1$
Then the smallest that the median value can be for a constructed election with seat share $S=\frac{1}{2}$ and vote share $V$ is a value approaching
$\frac{p+\frac{1}{2}}{2}=\frac{2V-1+\frac{1}{2}}{2}=V-\frac{1}{4}$
Thus, the infimal Mean-Median difference for $\frac{1}{2}\leq V\leq\frac{3}{4}$ is $-\frac{1}{4}$. Using symmetry, we get
[…]
Case 2: $S>\frac{1}{2}$, $\frac{1}{4}\leq V<\frac{3}{4}$
Firstly note that, since we assume that $S>\frac{1}{2}$, the median vote share must be at least as large as $\frac{1}{2}$. For $V\leq\frac{1}{2}$, we can achieve a median vote share of $\frac{1}{2}$ by constructing $S$ districts to have vote share $\frac{1}{2}$ (considered winning districts) and choosing the losing districts to have vote shares so that $\frac{1}{n}\sum_{i=1}^{n}V_{i}=V$, since we must have $S\leq 2V$. This would give vote share $V$, seat share $S$, and the smallest possible median of $\frac{1}{2}$.
For $\frac{1}{2}<V<\frac{3}{4}$, we can choose half of the districts to have vote share $\frac{1}{2}$ ($\ell$ of those considered to be losing districts, the rest winning) and then half to have vote share $V+\left(V-\frac{1}{2}\right)$. This would give vote share $V$, seat share $S$, and the smallest possible median of $\frac{1}{2}$.
Thus, we have so far shown that we can construct election data with minimum MM of
$\frac{1}{2}-V$
For the maximum possible median giving $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$, we would choose $\ell$ districts to have vote share 0, $\frac{n}{2}-\ell$ to have vote share $\frac{1}{2}$ (considered winning districts), and the remaining $\frac{n}{2}$ districts to have vote share $V_{max}$ that results in $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. This $V_{max}$ corresponds to the maximum median. We solve for $V_{max}$ below [^10] This argument glosses over the fact that the median value must be directly in the middle, but by choosing a large number of odd districts and using a limiting argument, we can approach the value stated here.:
$\frac{1}{n}\left(\frac{n}{2}V_{max}+\left(\frac{n}{2}-\ell\right)\frac{1}{2}\right)$ $=V$
$V_{max}=2V-S+\frac{1}{2}$
Of course, for $S<2V-\frac{1}{2}$, this $V_{max}$ is larger than 1, which cannot occur in a district’s vote share. In that case, the largest the median value can be is 1. Thus, for these $(V,S)$, we can construct corresponding election data having
$\frac{1}{2}-V$ $<MM<\min\left\{1-V,V-S+\frac{1}{2}\right\}$
Case 3: $S>\frac{1}{2}$, $\frac{3}{4}\leq V\leq 1$
For $V$ in this range, recall that $S\geq 2V-1$. Thus, we can construct election with half the districts to have vote share 1, $\ell$ to have vote share $\frac{1}{2}$, and the remaining districts to have vote shares resulting in $\frac{1}{n}\sum_{i=1}^{n}V_{i}=V$. This results in seat share $S$ and a maximum median value of 1.
For the minimum possible median giving $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$, we would choose $\ell$ districts to have vote share $\frac{1}{2}$ (corresponding to losing districts), $\frac{n}{2}$ to have vote share $1$ , and the remaining $\frac{n}{2}-\ell$ districts to have vote share $V_{min}$ that results in $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. This $V_{min}$ corresponds to the minimum median. We solve for $V_{min}$ below [^11] Again, we’re glossing over the fact that the median value must be directly in the middle, but by choosing a large number of odd districts and using a limiting argument, we can approach the value stated here.:
$\frac{1}{n}\left(\ell\cdot\frac{1}{2}+\left(\frac{n}{2}-\ell\right)V_{min}+\frac{n}{2}\right)$ $=V$
$\left(\frac{1}{2}-\left(1-S\right)\right)V_{min}$ $=V-\frac{1}{2}-\frac{\ell}{2n}=V-\frac{1}{2}-\frac{1}{2}\left(1-S\right)$
$V_{min}$ $=\frac{2V+S-2}{2S-1}$
Thus, we have shown the following bounds for the Mean-Median Difference:
$\frac{2V+S-2}{2S-1}-V$ $<MM<1-V$
Proof of Theorem 3.
For ease of exposition, we state the following in terms of the Partisan Bias, but note that the same examples will work for the Mean-Median Difference as well.
Suppose $S=\frac{n-\ell}{n}\geq\frac{1}{2}$ and vote shares $V_{1},V_{2},\dots,V_{n}$ give a Partisan Bias value of 0. In order to achieve the smallest possible vote share $V^{*}$ with turnout ratio $C$ corresponding to seat share $S$, we need to find $V_{i}$s which allow for the lowest possible vote shares in losing districts, and we need to weight those the most heavily. This corresponds to:
$V_{1}=V_{2}=\cdots=V_{\ell}$ $=0$
$V_{\ell}=V_{\ell+1}=\cdots=V_{n-\ell-1}$ $=\frac{1}{2}$
$V_{n-\ell}=V_{n-\ell+1}=\cdots=V_{n}$ $=1$
(Here districts with vote shares $\frac{1}{2}$ are winning).
The districts numbered 1 through $\ell$ will have $C$ times as much weight as the others, which implies they are each weighted $\frac{C}{n-\ell+C\ell}$, while districts numbered $\ell+1$ through $n$ are weighted $\frac{1}{n-\ell+C\ell}$. Now we can calculate the vote share $V^{*}$:
$V^{*}$ $=0\cdot\ell\cdot\frac{C}{n-\ell+C\ell}+(n-2\ell)\cdot\frac{1}{2}\cdot\frac{1}{n-\ell+C\ell}+\ell\cdot 1\cdot\frac{1}{n-\ell+C\ell}$
$=\frac{n}{2(n-\ell+C\ell)}$
Now we use the fact that $S=\frac{n-\ell}{n}$ to get
$V^{*}$ $=\frac{n}{2(nS+C(n-nS))}$
$=\frac{1}{2(S+C(1-S))}$
By adjusting the weight $C$, we can get any $V^{*}$ in between $\frac{1}{2(S+C(1-S))}$ and $\frac{1}{2}$.
Now we prove the upper bound by similar means. Suppose $S=\frac{n-\ell}{n}\geq\frac{1}{2}$ and vote shares $V_{1},V_{2},\dots,V_{n}$ give a Partisan Bias value of 0. In order to achieve the largest possible vote share $V^{*}$ with turnout ratio $C$ corresponding to seat share $S$, we need to find $V_{i}$s which allow for the largest number of vote shares at value 1, along with the largest possible vote shares in losing districts. Using the results from Corollary 1, we know this corresponds to:
$V_{1}=V_{2}=\cdots=V_{\ell}$ $=\frac{1}{2}$
$V_{\ell+1}=V_{\ell+2}=\cdots=V_{n/2}$ $=V$
$V_{n/2+1}=V_{n-\ell+1}=\cdots=V_{n}$ $=1$
where $V=\frac{2-S}{3-2S}$, which we get by solving the equation $S=\frac{3V-2}{2V-1}$ for $V$. (Here districts with vote shares $\frac{1}{2}$ are losing).
The districts numbered $n/2+1$ through $n$ will have $C$ times as much weight as the others, which implies they are each weighted $\frac{2C}{n(C+1)}$, while districts numbered $1$ through $n/2$ are weighted $\frac{2}{n(C+1)}$. Now we can calculate the vote share $V^{*}$:
$V^{*}$ $=\frac{1}{2}\cdot\ell\cdot\frac{2}{n(C+1)}+\left(\frac{n}{2}-\ell\right)\cdot\frac{2-S}{3-2S}\cdot\frac{2}{n(C+1)}+\frac{n}{2}\cdot\frac{2C}{n(C+1)}$
Again we use the fact that $S=\frac{n-\ell}{n}=1-\frac{\ell}{n}$ to get
$V^{*}$ $=\frac{1-S}{C+1}+\frac{2-S}{3-2S}\cdot\frac{1}{C+1}-\frac{2(1-S)(2-S)}{(C+1)(3-2S)}+\frac{C}{C+1}$
$=\frac{1+C(3-2S)}{(C+1)(3-2S)}$
By adjusting the weight $C$, we can get any $V^{*}$ in between $\frac{2-S}{3-2S}$ and $\frac{1+C(3-2S)}{(C+1)(3-2S)}$.
From Corollary 1, we already know that we can get any $V^{*}$ between $\frac{1}{2}$ and $\frac{2-S}{3-2S}$, so we are now done.
Appendix B Neutral ensemble Images
Figure 18: Neutral ensembles for Massachusett’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric. Declination value of 2 used for maps where Declination is undefined.
Figure 19: Neutral ensembles for Michigan’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 20: Neutral ensembles for Michigan’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 21: Neutral ensembles for Michigan’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 22: Neutral ensembles for Oregon’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 23: Neutral ensembles for Oregon’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 24: Neutral ensembles for Oregon’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 25: Neutral ensembles for Oklahoma’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric. Declination not included because it is not defined for a majority of maps.
Figure 26: Neutral ensembles for Oklahoma’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 27: Neutral ensembles for Oklahoma’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 28: Neutral ensembles for Pennsylvania’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 29: Neutral ensembles for Pennsylvania’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 30: Neutral ensembles for Texas’s State Senate districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Figure 31: Neutral ensembles for Texas’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Appendix C Short Burst Images
Figure 32: Short burst results for Massachusetts’s Congressional map and 2018 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for nearly all maps.
Figure 33: Short burst results for Massachusetts’s Congressional map and 2018 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for nearly all maps.
Figure 34: Short burst results for Massachusetts’s State Senate map and 2018 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for several maps.
Figure 35: Short burst results for Massachusetts’s State Senate map and 2018 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for several maps.
Figure 36: Short burst results for Michigan’s Congressional map and 2016 Presidential election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 37: Short burst results for Michigan’s Congressional map and 2016 Presidential election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 38: Short burst results for Michigan’s State Senate map and 2016 Presidential election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 39: Short burst results for Michigan’s State Senate map and 2016 Presidential election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 40: Short burst results for Michigan’s State House map and 2016 Presidential election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 41: Short burst results for Michigan’s State House map and 2016 Presidential election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 42: Short burst results for Oklahoma’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for a majority of the maps.
Figure 43: Short burst results for Oklahoma’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for a majority of the maps.
Figure 44: Short burst results for Oklahoma’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for a majority of the maps.
Figure 45: Short burst results for Oklahoma’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won. Declination not included because it is undefined for a majority of the maps.
Figure 46: Short burst results for Oklahoma’s State House map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 47: Short burst results for Oklahoma’s State House map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 48: Short burst results for Oregon’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 49: Short burst results for Oregon’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 50: Short burst results for Oregon’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 51: Short burst results for Oregon’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 52: Short burst results for Oregon’s State House map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 53: Short burst results for Oregon’s State House map and 2018 Gubenatorial election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 54: Short burst results for Pennsylvania’s Congressional map and 2016 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 55: Short burst results for Pennsylvania’s Congressional map and 2016 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 56: Short burst results for Pennsylvania’s State Senate map and 2016 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 57: Short burst results for Pennsylvania’s State Senate map and 2016 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 58: Short burst results for Pennsylvania’s State House map and 2016 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 59: Short burst results for Pennsylvania’s State House map and 2016 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 60: Short burst results for Texas’s Congressional map and 2014 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 61: Short burst results for Texas’s Congressional map and 2014 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 62: Short burst results for Texas’s State Senate map and 2014 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 63: Short burst results for Texas’s State Senate map and 2014 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 64: Short burst results for Texas’s State House map and 2014 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 65: Short burst results for Texas’s State House map and 2014 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.